OpenStudy (anonymous):

hat is the solution to the rational equation (3/2) - (1/x-4) = (-2/2x-8) ?

hero (hero):

Hint: \(\large\frac{3}{2} - \frac{1}{x-4} = -\frac{2}{2(x-4)}\)

hero (hero):

Cancel the 2/2 on the right side to get \(\large\frac{3}{2} - \frac{1}{x-4} = \frac{1}{x-4}\) Then add 1/(x-4) to both sides to get \(\large\frac{3}{2} = \frac{1}{x-4} + \frac{1}{x-4}\)

hero (hero):

You should be able to finish the rest from there

OpenStudy (anonymous):

i got 14/3

OpenStudy (anonymous):

???

hero (hero):

I got 16 in the numerator. Double check your work.

OpenStudy (anonymous):

then what would the final answer be , no solution right ?

hero (hero):

Why would it be no solution? x = 16/3 How is that no solution?

OpenStudy (anonymous):

this is what i have to pick from : No solution x = -4 x = 2 x = 4

hero (hero):

Oh...I made a mistake

hero (hero):

You should have gotten 3/2 = 0 after adding 1/(x-4) to both sides leaving no solution.

hero (hero):

3/2 = 0 cannot be true, therefore no solution.

OpenStudy (anonymous):

okay then .